Solve by the method of variation of of parameter y''+y=secx

1
Expert's answer
2022-02-10T13:50:37-0500

Solution:

Given "y^{\prime \prime}+y=\sec x" .

The auxilary equation is "r^{2}+1=0"

Then, "r^{2}=\pm i \Rightarrow r=-i, r=i ."

Thus, the general solution is, "y_{c}=c_{1} \cos x+c_{2} \sin x ."

Here, "y_{1}(x)=\cos x"  and "y_{2}(x)=\sin x" .

Then, "y_{1}^{\prime}(x)=-\sin x\ and\ y_{2}^{\prime}(x)=\cos x ."

Find the Wronskian of "y_{1}(x)=\cos x"  and "y_{2}(x)=\sin x ."

"\begin{aligned}

W &=\left|\begin{array}{cc}

\cos x & \sin x \\

-\sin x & \cos x

\end{array}\right| \\

&=\cos ^{2} x+\sin ^{2} x \\

&=1

\end{aligned}"

Here, "f(x)=\sec x ."

"\begin{aligned}

&u=-\int \frac{f(x) y_{2}(x)}{W(x)} d x \text { and } v=\int \frac{f(x) y_{1}(x)}{W(x)} d x \\

&u=-\int \frac{\sec x \sin x}{1} d x \quad \text { and } v=\int \frac{\sec x \cos x}{1} d x

\end{aligned}"

"\begin{array}{ll}

u=-\int \tan x d x & \text { and } v=\int 1 d x \\

u=\ln (\cos x) & \text { and } v=x

\end{array}"

Thus, the particular solution is,

"y_{p}=\cos x \ln (\cos x)+x \sin x"

Therefore, the general solution of the given differential equation is,

"\begin{aligned}

&y=y_{c}+y_{p} \\

&y=c_{1} \cos x+c_{2} \sin x+\cos x(\ln (\cos x))+x \sin x

\end{aligned}"